Two blocks each of mass 3 kg are connected by light cord and placed on rough horizontal surface when force if 20 newton is applied on the blocks an acceleration if .50m/s is produced in each block find the tension in the string and the frictional force which is same on the two blocks

Grade:11

1 Answers

Krishan Pareek
6 years ago
total external force =ma =(3+3)*(.5)= 3Nnow net force = 20-2f = 3.by solving this we get frictional force f= 8.5Nnow for the block which is in backward T-f= 3(0.50)= 1.5 therefore T = 1.5+f = 1.5+8.5 = 10 N